OpenStudy (anonymous):

factor the trinomial completely. if the trinomial contains a gcf (other than 1) factor out the gcf first. 2x^3-14x^2+20x

OpenStudy (anonymous):

=2x(x^2-7x+10) =2x(x^2-5x-2x+10) =2x[x(x-5)-2(x-5)] =2x(x-5)(x-2)